OpenStudy (anonymous):

My Calculus professor gave us this algebra question for extra credit. The first person in the class to solve the problem gets 5% added to their FINAL GRADE! The problem; g(x) = x^2 x =< 2 6-x x > 2 Find a single algebraic solution that satisfies both of these conditions. I hardly understand what he is asking us to do, let alone how to solve it. What I thought I heard him say was that when x =< 2, g(x) = x^2, when x > 2, g(x) = 6-x. Any help would be appreciated!

OpenStudy (anonymous):

Try plotting these two equation

OpenStudy (anonymous):

Ok, so there is a standard parabola and a line going through it intersecting at (2,4) and (-3,9).

OpenStudy (anonymous):

x^2=6-x x^2+x-6=0 (x+3)(x-2) this tell us that solution is x=-3,2 but that's not enough

OpenStudy (anonymous):

You might be waiting a while, hah. I was looking at the graph and that work and thinking that they were solutions. I am fairly terrible at intuitively working something out when I have no framework.

OpenStudy (anonymous):

Hold on , we are gonna get you that 5% EC

OpenStudy (anonymous):

Thanks so much! I'll work on this all night if I have to!

OpenStudy (anonymous):

Is it because x^2 =/= 6-x? For example, 1^2 =/= 6-1

OpenStudy (anonymous):

I am having second thought, what did you cover in your calc class?

OpenStudy (anonymous):

As far as I know this has nothing to do with calculus. In either case, today we covered some implicit differentiation, related rates, and continuity. He had said that only one student had ever solved this problem and that was with some help from the professor.

OpenStudy (anonymous):

I am now sure this is a limit problem

OpenStudy (anonymous):

Too bad you aren't a student of his or you could have made it two students. Hmm, maybe I will make it two students, but with help from you instead of him! :)

OpenStudy (anonymous):

Here is what I think, refer to graph: http://www.wolframalpha.com/input/?i=x^2%2C6-x&t=crmtb01 up to and including 2, we use x^2; from and not including 2, we use 6-x Let's evaluate both equation at 2 2^2=4 6-2=4 So I think solution is 4

OpenStudy (anonymous):

But that is the solution to the question what is the limit as x->2 for g(x).

OpenStudy (anonymous):

I will come back to this, but I am pretty sure this is right

OpenStudy (anonymous):

You're completely right, but as far as I know you did not answer the right question. He was not asking for the limit. He wants a single algebraic function, as opposed to its current form, which is piecewise.

OpenStudy (anonymous):

The way he worded it was that you would get an answer like; g(x) = x^2+4x-6 (arbitrary) When x =< 2 g(x) = x^2 when x > 2 g(x) = 6-x

OpenStudy (anonymous):

I will look at this tonight

OpenStudy (anonymous):

Alright, great, thanks for your help and any other help you can provide!

myininaya (myininaya):

is he just looking for the roots? (x<=2) : x^2=0 when x=0 (x>2): 6-x=0 when x=6 the graph touches the x-axis when x=0 and x=6

OpenStudy (anonymous):

no, one function that looks like this piecewise function

OpenStudy (anonymous):

Yeah, what Imranmeah91 said. I really have no idea where to begin. Such a weird problem, but I guess that is why only one of his students have solved it.

myininaya (myininaya):

well i'm kindof thinking we are going to see absolute value in this y=|x-2|+4 is close looking maybe this could be a starting point

myininaya (myininaya):

y=|x-2|+4 works great for the left hand side of the graph but we need to work on the curve of the right hand side

myininaya (myininaya):

you know who i think i would be awesome to solve this problem that guy that calls himself mathteacher he likes to use technology to solve problems hes awesome at it

OpenStudy (anonymous):

Very nice, Myininaya! That is a great starting point. How did you figure that out, just intuition? What boggles my mind is how to get the other half of the equation without DRASTICALLY altering the first half.

myininaya (myininaya):

yea i just seen it had sort of the same graph of |x| but i'm kindof stumped i don't know if we can do this lol

myininaya (myininaya):

he wants to make that piecewise function into one function

OpenStudy (anonymous):

the way i read it is \[ g(x) = \left\{ \begin{array}{lr} x^2 & : x \leq 2\\ 6-x & : x >2 \end{array} \right. \] a piecewise function that cannot be written as a single algebraic expression so don't bother trying. do you have the question exactly as written?

OpenStudy (anonymous):

it is clearly not one expression which is why it is piecewise. maybe the problem really is something else

myininaya (myininaya):

i was kindof thinking that

OpenStudy (anonymous):

maybe it is two inequalities to solve

myininaya (myininaya):

he said only one student has been able to solve it in the past so i guess whatever it is its not really easy

OpenStudy (anonymous):

maybe it is a joke. this has nothing to do with calc in any case

myininaya (myininaya):

and that student had the help of a professor

OpenStudy (anonymous):

one is a line. the other is a parabola what do you want, a miracle?

myininaya (myininaya):

yes

OpenStudy (anonymous):

i will give it some thought while i go have a beer.

myininaya (myininaya):

i think you are right though satellite this is impossible

OpenStudy (anonymous):

haha, Ill take a miracle. Yeah, I'm not sure what to tell you. Unfortunately he did not write the exact question down, only wrote that piecewise function and told us to come up with an a single algebraic function.

OpenStudy (anonymous):

I am sending the professor an email to ensure that I have the right understanding of what answer he is looking for.

OpenStudy (anonymous):

Quick update; I sent my professor an email ensuring that I understood the question and to show him the progress that was made so far from this thread. Here is his response; "You do seem to have the idea, but are you sure that [using abs(x) for absolute value of x] that your expression, abs(x-2) + 4, is algebraic? Can you really write it solely in terms of add, subtract, multiply, divide, power, and root? Because that’s what “algebraic” means. Now in fact that function you wrote is algebraic, but maybe as a first challenge you should actually write such a formula for it." So there must be some way to write the function in terms of ASMDPR and this is the path I am going to pursue now, which will hopefully give me a better understanding of how to alter the function y = |x-2| + 4

OpenStudy (anonymous):

I am gonna play with Mathematica to see if I can generate such function

OpenStudy (anonymous):

Wow, thought I might try to do the same until I realized it has a $300 price tag!

myininaya (myininaya):

ok joe we have a piecewise function g(x)= x^2 for x>=2 and 6-x for x<2 the goal is to write using one algebra function

myininaya (myininaya):

i dont see how it is possible but it seems the professor knows it is

OpenStudy (anonymous):

I sure hope this professor is not trying to make a point

OpenStudy (anonymous):

I don't think he would respond to my email in such a fashion if he was just trying to trick us. Plus, this is a calculus class. If he was going to trick us for extra credit I doubt it would be an algebra problem.

myininaya (myininaya):

the closer thing i could come up with is y=|x-2|+4

OpenStudy (anonymous):

yeah, i drew a rough picture of it but nothing is coming to mind immediately. Do we know what the student has learned right before this was asked? like what the topic is over or something?

OpenStudy (anonymous):

Yeah, we were covering related rates, continuity, and implicit differentiation. He used this to prove some point about continuity (I believe), but the extra credit question had nothing to do with what we were learning. He did say that only one student had ever solved this problem and that was with some help from the professor himself.

myininaya (myininaya):

or maybe it would be an algebra problem since we are supposedly to know algebra

OpenStudy (anonymous):

What is the equation for that line, imranmeah91?

myininaya (myininaya):

i believe no pattern exist for the number sequence 6,5,4,9,16,... does it? hey joe can you come up with an expression like you did that one time for that one pattern that no one else could get

OpenStudy (anonymous):

How about this?

OpenStudy (anonymous):

i can create a formula for that sequence, but i dont know if that would help, i dont have that much control over the tail of the sequence.

OpenStudy (anonymous):

i wish there what a natural sequence that what like, u(t) = 0 for x <0, and 1 for x>= 0, that would solve this in a heart beat.

OpenStudy (anonymous):

Did you see my graph?

myininaya (myininaya):

:( ok don't waste your time then

OpenStudy (anonymous):

how did you get that imranmeah?

OpenStudy (anonymous):

Yeah imran, what is the formula for that graph? Looks damn close.

OpenStudy (anonymous):

He used mathematica I believe

OpenStudy (anonymous):

I used mathematica, it is not exact but you may be able to adjust

OpenStudy (anonymous):

it needs to be shifted to the right some.

OpenStudy (anonymous):

New graph and equation hold on

OpenStudy (anonymous):

im posting a (crude) pic of what the function looks like.

OpenStudy (anonymous):

just follow the solid lines.

OpenStudy (anonymous):

\[4-\sqrt[3]{(2-x)^2} (x+6)^{2/3}\]

OpenStudy (anonymous):

I think it need to be adjusted from here

OpenStudy (anonymous):

the best thing i can think of here is to use an indicator funtion

myininaya (myininaya):

oops i have been doing it for x^2,x>=2 and 6-x, x<2

OpenStudy (anonymous):

let g(x)=x^2 * I where I=1 when x<=2 and I=(6-x)/x^2 when x>2...but it seems simple

OpenStudy (anonymous):

Imranmeah, after putting that function into my graphing calculator I one problem in that anything anything < -6 is not a part of the parabola. When x = -8 y should be 64, but with that function it is -3.36. Damn good start though.

OpenStudy (anonymous):

i think i might be on to something. Check this out: http://www.wolframalpha.com/input/?i=%281%2F%284-2x%29%29%5B |2-x|%2B%282-x%29%5Dx^2%2B%281%2F%282x-4%29%29%5B|x-2|%2B%28x-2%29%5D%286-x%29

OpenStudy (anonymous):

That's perfect

OpenStudy (anonymous):

this was my idea: I want to make some function like: \[h(x) = ax^2+b(6-x)\] where a = 1, b = 0 when x is less than 2, and a = 0, b = 1 when x is greater than 2.

OpenStudy (anonymous):

so the trouble comes in finding some functions like that, but what i ended up using at first is: a =|2 - x| + (2 - x) b = |x - 2| + (x - 2) because when x is less than 2, b will equal 0, and when x is greater than 2, a will equal 0.

OpenStudy (anonymous):

thats the same idea as trying to make an indicator function, where ur indicators a nd b take on different values at different values of x

OpenStudy (anonymous):

My god, Joe, brilliant. I have yet to plug the numbers in to check, but this looks great so far.

OpenStudy (anonymous):

but their values arent 1 when their respective conditions are satisfied: if x is less than 2, a becomes: 2(2 - x) = 4 - 2x and if x is greater than 2, b becomes: 2(x - 2) = 2x-4 So i just divide by those values to make the constant come out to 1 thus: \[a = \frac{1}{4-2x}(|2-x|+(2-x)), b = \frac{1}{2x-4}(|x-2|+(x-2))\] and the equation you are looking for is: \[g(x) = ax^2+b(6-x)\]

OpenStudy (anonymous):

Joe, when plugging the number 2 into the formula, which should give an answer of four, came out indeterminate. Any idea why?

OpenStudy (anonymous):

because of my fractions >.< let me see if i can think of some way to take care of that.

OpenStudy (anonymous):

1/(4-2x) = 1/0 when x=2

OpenStudy (anonymous):

i got nothing, i dont see anyway to avoid the discontinuity at x = 2 =/ other than that, does it seem good? anyone have any objections to that?

OpenStudy (anonymous):

I think it is a good as it is gonna get

OpenStudy (anonymous):

wow joe that is amazing!

myininaya (myininaya):

that is my cup of joe i dont share

OpenStudy (anonymous):

but i so have one small (very very small) criticism namely that you are using absolute value

OpenStudy (anonymous):

Yeah, that really was amazing Joe. Absolutely brilliant. I think I might have seen you on Good Will Hunting ;)

OpenStudy (anonymous):

which itself is piecewise.

myininaya (myininaya):

shh satellite

OpenStudy (anonymous):

^ Although my professor did note that it was acceptable to use it.

OpenStudy (anonymous):

sorry i go back to the dungeon now

OpenStudy (anonymous):

thanks :) i got the idea when myininaya posted that abs function. so thank you myininaya! and now ive learned something new, yay :) thats why i like this site lol.

myininaya (myininaya):

gj joe :)

OpenStudy (anonymous):

i learn something here every day (almost)